OpenStudy (auw7):

How do i solve the rest of this logarithmic equation log(subscript16)32=x+2. So far I have 32=16^(x+2) x+2=log(subscript16)32. Don't know how to solve from here can someone please help me?

OpenStudy (auw7):

I think x=-0.75 now I am checking my work

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle16^{x+2}=32 }\) \(\color{#000000 }{ \displaystyle \log_{16}16^{x+2}=\log_{16}32 }\) \(\color{#000000 }{ \displaystyle (x+2)\log_{16}16=\log_{16}32 }\) \(\color{#000000 }{ \displaystyle (x+2)\cdot 1=\log_{16}32 }\) \(\color{#000000 }{ \displaystyle x+2=\log_{16}32 }\)

OpenStudy (solomonzelman):

I am just confirming that your result for converting into logarithm is right

OpenStudy (auw7):

Really? Thanks!

OpenStudy (solomonzelman):

you can further simplify it by applying the rules \(\color{#000000 }{ \displaystyle x+2=\log_{16}32 }\) \(\color{#000000 }{ \displaystyle x+2=\log_{16}(16\cdot 2) }\) \(\color{#000000 }{ \displaystyle x+2=\log_{16}(16) +\log_{16}(2) }\) \(\color{#000000 }{ \displaystyle x+2=1 +\log_{16}(2) }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x=-1+\log_{16}(2) }\)

OpenStudy (solomonzelman):

and then use the fact that \(2^4=16\quad \Longrightarrow \sqrt[4]{16}\quad \Longrightarrow {16}^{1/4}\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x=-1+\log_{16}(2) }\) \(\color{#000000 }{ \displaystyle x=-1+\log_{16}(16^{1/4}) }\)

OpenStudy (solomonzelman):

so your answer is right, because after taking the exponent out you get exactly that

OpenStudy (auw7):

I understand that more

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x=-1+\log_{16}(16^{1/4}) }\) \(\color{#000000 }{ \displaystyle x=-1+(1/4)\log_{16}(16) }\) \(\color{#000000 }{ \displaystyle x=-1+(1/4)\cdot 1}\) \(\color{#000000 }{ \displaystyle x=-1+(1/4)}\) \(\color{#000000 }{ \displaystyle x=-1+0.25=-0.75 }\)

OpenStudy (solomonzelman):

good job\( : \)

OpenStudy (auw7):

Thank you!

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

\[16^{x+2}=32\] \[\left( 2^4 \right)^{x+2}=2^5\] \[2^{4\left( x+2 \right)}=2^5,4\left( x+2 \right)=5,4x=5-8=-3,x=-\frac{ 3 }{ 4 }\]